# Sustainable Self-evolution Adversarial Training

### v1

@@ 4.3. Ablation Study
Visualization experiments of clean samples representations. We additionally visualize the representations of clean examples with CAD baseline and SSEAT. As shown in Fig. 3, we can see that compared with the CAD baseline, SSEAT exhibits more distinct distances between classes in feature representations on clean samples, with closer intra-class distances, which enhances its discriminative ability. For example, in SSEAT, the distribution of red points representing Class 0 is more clustered, whereas in the CAD baseline, the distribution is more dispersed. This observation indicates that our SSEAT method obtains more discriminative representations of clean examples compared with the CAD baseline. These results indicate that our SSEAT can maintain high accuracy on clean samples.
@@ 5. Conclusion
In the real world, data is often subject to noise and disturbances, necessitating models with strong robustness and adaptability. To realize the adaptive defense ability of deep models with adversarial training when facing the continuously generated diverse new adversarial samples, we proposed a novel and task-driven Sustainable Self-evolution Adversarial Training (SSEAT) method. The SSEAT framework we propose aims to achieve autonomous model evolution and defense against various adversarial attacks, enabling models to effectively adapt to complex and changing environments, thereby enhancing their reliability and stability in multi-media practical applications. Inspired by the continue learning, the SSEAT framework can continuously learn new kinds of adversarial examples in each training stage, and realize the consolidation of old knowledge through the data rehearsal strategy of high-quality data selection. At the same time, a knowledge distillation strategy is used to further maintain model classification accuracy on clean samples when facing multiple adversarial attacks. We have verified the SSEAT model efficacy over multiple continue defense setting orders, and the ablation experiments show the effectiveness of the components in the SSEAT method.

### v2

@@ 5. Conclusion
To realize the adaptive defense ability of deep models with adversarial training when facing the continuously generated diverse new adversarial samples, we proposed a novel and task-driven Sustainable Self-evolution Adversarial Training (SSEAT) method. Inspired by the continue learning, the SSEAT framework can continuously learn new kinds of adversarial examples in each training stage, and realize the consolidation of old knowledge through the data rehearsal strategy of high-quality data selection. At the same time, a knowledge distillation strategy is used to further maintain model classification accuracy on clean samples. We have verified the SSEAT model efficacy over multiple continue defense setting orders, and the ablation experiments show the effectiveness of the components in the SSEAT method.